in a chocolate factory it took 20 workers 6 days to fil an order for chocolate valentine candy. How long would it take if there were 30 workers?
Answer by ankor@dixie-net.com(12704)

n a chocolate factory it took 20 workers 6 days to fil an order for chocolate valentine candy. How long would it take if there were 30 workers?
:
Represent the completed job in "man-days"
In this problem:
20 * 6 = 120 man-days
:
How many days to complete the same job with 30 men?
30*d = 120
d = 120/30
d = 4 days to do the same job with 30 guys.